tom is 4 times as old as todd. in 5 years toms age will be 1 year less than 3 times as old as todd. how old is tom?
Answer by dabanfield(803)

tom is 4 times as old as todd. in 5 years toms age will be 1 year less than 3 times as old as todd. how old is tom?
Let x be Todd's age and y Tom's. Then we have:
1.) y = 4*x
In 5 years their ages will be x+5 and y+5 so we have:
2.) y+5 = 3*(x+5)-1
We can substitute 4x for y in 2.) and solve for x:
4x+5 = 3*(x+5)-1
Once you have x you can calculate y = 4*x.